OpenStudy (anonymous):

MEDAL AND FAN The contents of a cell in a spreadsheet are shown below. 7.9709E+7 What number do the contents of the cell represent? 79,709,000 0.0000000797 797,090,000,000 0.0000007971

OpenStudy (drigobri):

Simple.. E+Number is = x10^(number) For instance, 3.25E+3 is 3.25x10³ = 3.25 x 1000 = 3250 On your case 7.9709E+7 = 7.9709 x 10^7 = 7.9709 x 10000000 = 79 709 000

OpenStudy (anonymous):

ok thank you